Jimmy is about to take part in the high dive. He jumps up off the high board at an angle of 23 degrees with an initial velocity of 2.5m/s. How far does Jimmy travel before he hits the water (horizontal distance). Consider air resistance


1
Expert's answer
2021-04-12T07:02:20-0400

horizontal distance:

"x=v_xt"

"v_x=v_0cos\theta"

"v_0" is initial velocity.

"\theta=23\degree"

vertical distance:

"y=v_0tsin\theta-gt^2/2"

When Jimmy hits the water:

"y=0"

"v_0sin\theta=gt/2"

"t=2v_0sin\theta/g=2\cdot2.5sin23\degree/9.8=0.2\ s"

"x=2.5\cdot0.2cos23\degree=0.46\ m"